OpenStudy (anonymous):

The Everton college store paid $1813 for an order of 45 calculators. The store paid $10 for each scientific calculator. The others, all graphing calculators cost the store $57 each. How many of each type of calculator was ordered?

OpenStudy (anonymous):

let the number of scientific calculators be x and number of graphing calculators be y x+y=45 10x+57y=1813 solve for x and y

OpenStudy (anonymous):

how do I solve it?

OpenStudy (anonymous):

x+y=45 ------------(1) 10x+57y=1813--------(2) multiply 1st equation by 10 10(x+y=45)------->10x+10y=450 subtract 1st equation from 2nd equation 10x+57y=1813 -10x-10y=-450 --------------- 47y=1363 y=1363/47 y=29 now plug y value into one of the equation x+29=45 x=45-29 x=16 therefore,number of scientific calculators=16 and number of graphing calculators=29

OpenStudy (anonymous):

OOOHHH okay, thank you :D